# Agentic Large Language Models, a survey

### v2

@@ 4.4.2 Use Cases: Emergent Behavior and Training Data
New Training Data
A final use case of this third part of the taxonomy is that new training data is generated by the interacting agents. Traditionally, LLMs are trained on a large static corpus of language data, that is taken from the internet, and ultimately based on human actions, using self-supervised learning methods. As illustrated by the cycle in Figure 1, interacting agentic LLMs enable self-learning, in the style of reinforcement learning. Reinforcement learning is used increasingly in LLM training, for example to train reasoning models by OpenAI (Huang et al., 2024c; Wu et al., 2024) and DeepSeek (Guo et al., 2025).
In reinforcement learning, agents choose their own actions in the world, and are not limited by a pre-existing dataset. In principle, they can learn the full complexity of the world, including the effects of their own actions. A challenge in reinforcement learning is the instability caused by feedback loops (Plaat, 2022). Past reinforcement learning successes have achieved stable training through diverse exploration and low learning rates, requiring large computational efforts (Silver et al., 2016; Vinyals et al., 2019; Brown and Sandholm, 2019). Open-ended and open-world multi-agent simulation may provide an alternative way to create the necessary diversity for stable convergence.
5 General Discussion and Research Agenda
@@
5.1 Research Agenda for Agentic LLM
@@
Such feedback loops are common in reinforcement learning, where agents act and receive feedback from their environment. In games, a self-learning loop can be created (Plaat, 2022). In AlphaGo Zero this approach yielded good results, although at the cost of careful tuning of hyperparameters and algorithms, to ensure sustained convergence of the learning process (Silver et al., 2017). Similar results were achieved in other challenging games, such as StarCraft (Vinyals et al., 2019), Stratego (Perolat et al., 2022), DOTA 2 (Berner et al., 2019), Diplomacy (Bakhtin et al., 2022), and Poker (Brown and Sandholm, 2019).
@@
The approach that worked well in games of strategy is now also successfully used in robotics, in the creation of Vision-Language-Action models (Black et al., 2024; Brohan et al., 2023). VLAs that are trained on self-generated action sequences show zero-shot generalization results in domestic tasks (kitchen tasks, folding laundry) that had not been achieved by other machine learning methods. Recently, reasoning models—such as DeepSeek (Guo et al., 2025) and Kimi (Du et al., 2025)—are also being trained with reinforcement learning. Other uses of agents for finetuning are reported by Subramaniam et al. (2025).
Reuse of inference time results for finetuning and pretraining closes the learning loop (see Figure 1), and is the first item for the agenda for further research. It is interesting to see how the reinforcement learning methods that worked well for games of strategy are being translated to work in LLMs that act in the real world.
@@
Faithfulness for Chain of Thought is studied by Lyu et al. (2023); Lanham et al. (2023); Turpin et al. (2024). Mechanistic interpretability can provide ways to look inside the LLM, to better understand if the model follows the reasoning steps that we expect it to take (Nanda et al., 2023; Bereska and Gavves, 2024; Ferrando et al., 2024). The conditions that influence stability of emergent behavior (cooperation, fairness, trust) may be studied further.
@@
Studies of emergent behavior need realistic agent behavior, and we expect more research to be performed to improve agent behavior, for example by closely modeling human behavior in generative agents (Park et al., 2024). Some behavior patterns in multi-agent simulations only emerge at scale, as studies with specialized agent infrastructures have shown (Park et al., 2023; Yang et al., 2024e; Wu et al., 2023). However, the number of LLM agents that can be simulated reliably is often limited. Although open-ended simulation show improved scalability, we believe that more research into scaling of simulations with complex agents is necessary.
Related to the challenge of scale is the cost of training LLMs. Pretraining and finetuning an LLM is expensive. Knowledge distillation is a popular method to extract essential knowledge and behavior from a large model into a small model, at lower computational cost (Xu et al., 2024b). Experiments have shown that reasoning steps can be distilled from large to smaller language models (Gu et al., 2023; Li et al., 2023b; Muennighoff et al., 2025). Knowledge distillation in LLM agents in an important item for our research agenda.
@@
Self reflection mechanisms are used in advanced prompt-improvement algorithms. Hand-writing external prompt management algorithms may be error prone and brittle. An alternative is to let the LLM perform the self reflection and step-by-step management internally, as in the original Chain of Thought.
DeepSeek R1 (Guo et al., 2025) is a reasoning model that is trained (finetuned) by the GRPO reinforcement learning method (Shao et al., 2024). The model is trained on its own reasoning results, and was found to self-reflectively reason over its own results, identifying effective reasoning patterns implicitly. Schultz et al. (2024) train a model on search sequences (Gandhi et al., 2024) in games such as chess, and VLAs are trained on action sequences (Kim et al., 2024). These works shows that, in addition to implicit step-by-step reasoning, implicit search is possible. An open question is whether LLMs can perform self reflection internally.
By adding external state to an LLM, we enable reasoning and a form of self reflection, which is a rudimentary form of metacognition (thinking about thinking). LLMs that reflect on their own behavior raise visions of true artificial intelligence. If LLMs can self-reflect, can they exhibit metacognition (Wang and Zhao, 2023; Didolkar et al., 2024)? Self reflection by LLMs is another item for the research agenda.
@@
Self reflective methods are being used to create agents to perform scientific discovery (Eger et al., 2025). How these agents will influence, and possibly improve, the process of scientific discovery is an exciting area of research.
@@
Safety is an important issue in LLMs that act in the world. The problem is studied, but far from being solved (Brunke et al., 2022; Andriushchenko et al., 2024; Samvelyan et al., 2024). Actions by assistants and robots in the real world have real world consequences. When a financial trading assistant hallucinates, or when a self driving robot makes a wrong inference, questions on responsibility and liability should be addressed. More legal and ethical questions arise, for example, on privacy and fairness, and, possibly, concerning the rights of algorithmic entities (Harris and Anthis, 2021).
The application areas for the assistants in this survey—shopping, medical diagnosis, finance—are narrow. The narrower the application domain, the better the answers. Clearly, many safety, ethics and trust issues will have to be addressed before the full breadth of the possibilities of agentic LLMs can be enjoyed. Safety will continue to be an important topic on the research agenda of agentic LLMs.
@@
5.2 Conclusion
There is a large research activity on agentic LLMs. Already, robots show impressive results, and so do assistants in medical diagnosis, financial market advising, and scientific research. Work processes in these—and other—fields may well be affected by agentic LLM assistants in the near future.
@@ 5.2 Conclusion
The agentic paradigm enables two elements of machine learning that are new for LLMs. In reinforcement learning, agents self reflect and choose their own actions, and learn from the feedback of the world in which they operate. No dataset is needed, nor is learning limited by the complexity of a dataset, it is only limited by the complexity of the world around the agent. The agent paradigm creates a more challenging training setting, allowing agentic LLMs to keep improving themselves.
@@ Acknowledgements
We thank Joost Broekens, Suzan Verberne, Annie Wong, Thomas Bäck, Zhaochun Ren, Thomas Moerland, Michiel van der Meer, Bram Renting, Frank van Harmelen, Tessa Verhoef, and Rob van Nieuwpoort for extensive and fruitful discussions.

### v3

@@ 4.4.3. Emergent Collective Behavior
4.4.4. New Training Data
A final use case of this third part of the taxonomy is that new training data is generated by the interacting agents. Traditionally, LLMs are trained on a large static corpus of language data, that is taken from the internet, and ultimately based on human actions, using self-supervised learning methods. As illustrated by the cycle in Figure 1, interacting agentic LLMs enable self-learning, in the style of reinforcement learning. Reinforcement learning is used increasingly in LLM training, for example to train reasoning models by OpenAI (huang2024o1; wu2024comparative), and DeepSeek (guo2025deepseek). New reinforcement learning methods such as GRPO (shao2024deepseekmath) and RLVR (lambert2024tulu) already allow inference-time chains of thoughts to be used for finetuning.
In reinforcement learning, agents choose their own actions in the world, and are not limited by a pre-existing dataset. In principle, they can learn the full complexity of the world, including the effects of their own actions. A challenge in reinforcement learning is the instability caused by feedback loops. Past reinforcement learning successes have achieved stable training through diverse exploration and low learning rates, requiring large computational efforts (silver2016mastering; vinyals2019grandmaster; brown2019superhuman). Open-ended and open-world multi-agent simulation may provide an alternative way to create the necessary diversity for stable convergence.
5. General Discussion and Research Agenda
@@
5.1. Research Agenda for Agentic LLM
@@
Such feedback loops are common in reinforcement learning, where agents act and receive feedback from their environment. In games, a self-learning loop can be created (plaat2022deep). In AlphaGo Zero this approach yielded good results, although at the cost of careful tuning of hyperparameters and algorithms, to ensure sustained convergence of the learning process (silver2017mastering). Similar results were achieved in other challenging games, such as StarCraft (vinyals2019grandmaster), Stratego (perolat2022mastering), DOTA 2 (berner2019dota), Diplomacy (meta2022human), and Poker (brown2019superhuman).
@@
The approach that worked well in games of strategy is now also successfully used in robotics, in the creation of Vision-Language-Action models (black2024pi_0; brohan2023rt). VLAs that are trained on self-generated action sequences show zero-shot generalization results in domestic tasks (kitchen tasks, folding laundry) that had not been achieved by other machine learning methods. Recently, reasoning models—such as DeepSeek (guo2025deepseek) and Kimi (team2025kimi)—are also being trained with reinforcement learning. Popular methods for finetuning for mathematics and coding tasks are GRPO (shao2024deepseekmath) and RLVR (lambert2024tulu). Other uses of agents for finetuning are reported by subramaniam2025multiagent.
Reuse of inference time results for finetuning and pretraining closes the learning loop (see Figure 1), and is the first item for the agenda for further research. It is interesting to see how the reinforcement learning methods that worked well for games of strategy and certain finetuning tasks, are being translated to work in LLMs that act in the real world.
@@
Faithfulness for Chain of Thought is studied by lyu2023faithful; lanham2023measuring; turpin2024language. Mechanistic interpretability can provide ways to look inside the LLM, to better understand if the model follows the reasoning steps that we expect it to take (nanda2023progress; bereska2024mechanistic; ferrando2024primer; chen2025does). The conditions that influence stability of emergent behavior (cooperation, fairness, trust) may be studied further.
@@
Studies of emergent behavior need realistic agent behavior, and we expect more research to be performed to improve agent behavior, for example by closely modeling human behavior in generative agents (park2024generative). Some behavior patterns in multi-agent simulations only emerge at scale, as studies with specialized agent infrastructures have shown (park2023generative; yang2024oasis; wu2023autogen). However, the number of LLM agents that can be simulated reliably is often limited. Although open-ended simulation show improved scalability, we believe that more research into scaling of simulations with complex agents is necessary.
Related to the challenge of scale is the cost of training LLMs. Pretraining and finetuning an LLM is expensive. Knowledge distillation is a popular method to extract essential knowledge and behavior from a large model into a small model, at lower computational cost (xu2024survey). Experiments have shown that reasoning steps can be distilled from large to smaller language models (gu2023minillm; li2023symbolic; muennighoff2025s1). Knowledge distillation in LLM agents in an important item for our research agenda.
@@
Self reflection mechanisms are used in advanced prompt-improvement algorithms. Hand-writing external prompt management algorithms may be error prone and brittle. An alternative is to let the LLM perform the self reflection and step-by-step management internally, as in the original Chain of Thought (implicit reasoning (li2025implicit)).
DeepSeek R1 (guo2025deepseek) is a reasoning model that is trained (finetuned) by the GRPO reinforcement learning method (shao2024deepseekmath). The model is trained on its own reasoning results, and was found to self-reflectively reason over its own results, identifying effective reasoning patterns implicitly. schultz2024mastering train a model on search sequences (gandhi2024stream) in games such as chess, and VLAs are trained on action sequences (kim2024openvla). These works shows that, in addition to implicit step-by-step reasoning, implicit search is possible. An open question is whether LLMs can perform self reflection internally.
By adding external state to an LLM, we enable reasoning and a form of self reflection, which is a rudimentary form of metacognition (thinking about thinking). LLMs that reflect on their own behavior raise visions of true artificial intelligence. If LLMs can self-reflect, can they exhibit metacognition (wang2023metacognitive; didolkar2024metacognitive)? Self reflection by LLMs is another item for the research agenda.
@@
Self reflective methods are being used to create agents to perform scientific discovery (eger2025transforming). How these agents will influence, and possibly improve, the process of scientific discovery is an exciting area of research.
@@
Safety is a crucial issue in LLMs that act in the world. The problem is studied, but far from being solved (brunke2022safe; andriushchenko2024agentharm; samvelyan2024rainbow). Actions by assistants and robots in the real world have real world consequences. When a financial trading assistant hallucinates, or when a self driving robot makes a wrong inference, questions on responsibility and liability should be addressed. More legal and ethical questions arise, for example, on privacy and fairness, and, possibly, concerning the rights of algorithmic entities (harris2021moral; bengio2025illusions).
The application areas for the assistants in this survey—shopping, medical diagnosis, finance—are narrow. The narrower the application domain, the better the answers.
Ensuring the safety of agentic LLMs requires moving beyond prompt-based defenses toward integrated, multi-layer safeguards. Key directions include embedding explicit safety constraints within the agent’s reasoning and planning pipeline and employing continual adversarial training and automated red-teaming to enhance robustness against manipulation. Further priorities are developing mechanisms for self-regulation and risk awareness, enabling agents to detect and avoid unsafe actions, and establishing rigorous, standardized safety benchmarks such as AgentHarm (andriushchenko2024agentharm). Together, these measures outline a roadmap for accountable and trustworthy deployment of agentic LLMs in high-stakes domains.
Clearly, many safety, ethics and trust issues will have to be addressed before the full breadth of the possibilities of agentic LLMs can be enjoyed. Safety and ethics will become important topics on the research agenda of agentic LLMs, deserving their own surveys and books (gan2024navigating; jiao2025navigating; raza2025responsible).
@@
5.2. Conclusion
There is a large research activity on agentic LLMs. Already, robots show impressive generalization results, and so do assistants in medical diagnosis, financial market advising, and scientific research. Work processes in these—and other—fields may well be affected by agentic LLM assistants in the near future.
@@ 5.2. Conclusion
The agentic paradigm enables two elements of machine learning that are new for LLMs. First, in reinforcement learning, agents self reflect and choose their own actions, and learn from the feedback of the world in which they operate. Second, no dataset is needed, nor is learning limited by the complexity of a dataset, it is only limited by the complexity of the world around the agent. The agent paradigm creates a more challenging training setting, allowing agentic LLMs to keep improving themselves.
@@ Acknowledgements
We thank Joost Broekens, Suzan Verberne, Annie Wong, Thomas Bäck, Zhaochun Ren, Thomas Moerland, Michiel van der Meer, Bram Renting, Frank van Harmelen, Tessa Verhoef, and Rob van Nieuwpoort for extensive and fruitful discussions. We thank the anonymous reviewers for valuable suggestions that have improved the article.
6. Reproducibility Checklist for JAIR
Select the answers that apply to your research – one per item.
All articles:
All claims investigated in this work are clearly stated. [yes]
Clear explanations are given how the work reported substantiates the claims. [yes]
Limitations or technical assumptions are stated clearly and explicitly. [yes]
Conceptual outlines and/or pseudo-code descriptions of the AI methods introduced in this work are provided, and important implementation details are discussed. [yes]
Motivation is provided for all design choices, including algorithms, implementation choices, parameters, data sets and experimental protocols beyond metrics. [yes]
Articles containing theoretical contributions:
Does this paper make theoretical contributions? [no]
If yes, please complete the list below.
All assumptions and restrictions are stated clearly and formally. [yes/partially/no]
All novel claims are stated formally (e.g., in theorem statements). [yes/partially/no]
Proofs of all non-trivial claims are provided in sufficient detail to permit verification by readers with a reasonable degree of expertise (e.g., that expected from a PhD candidate in the same area of AI). [yes/partially/no]
Complex formalism, such as definitions or proofs, is motivated and explained clearly. [yes/partially/no]
The use of mathematical notation and formalism serves the purpose of enhancing clarity and precision; gratuitous use of mathematical formalism (i.e., use that does not enhance clarity or precision) is avoided. [yes/partially/no]
Appropriate citations are given for all non-trivial theoretical tools and techniques. [yes/partially/no]
Articles reporting on computational experiments:
Does this paper include computational experiments? [no]
If yes, please complete the list below.
All source code required for conducting experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for source code readability and documentation as well as for long-term accessibility. [yes/partially/no]
The source code comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
The source code comes with a license that allows free usage for research purposes in general. [yes/partially/no]
Raw, unaggregated data from all experiments is included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility. [yes/partially/no]
The unaggregated data comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
The unaggregated data comes with a license that allows free usage for research purposes in general. [yes/partially/no]
If an algorithm depends on randomness, then the method used for generating random numbers and for setting seeds is described in a way sufficient to allow replication of results. [yes/partially/no/NA]
The execution environment for experiments, the computing infrastructure (hardware and software) used for running them, is described, including GPU/CPU makes and models; amount of memory (cache and RAM); make and version of operating system; names and versions of relevant software libraries and frameworks. [yes/partially/no]
The evaluation metrics used in experiments are clearly explained and their choice is explicitly motivated. [yes/partially/no]
The number of algorithm runs used to compute each result is reported. [yes/no]
Reported results have not been “cherry-picked” by silently ignoring unsuccessful or unsatisfactory experiments. [yes/partially/no]
Analysis of results goes beyond single-dimensional summaries of performance (e.g., average, median) to include measures of variation, confidence, or other distributional information. [yes/no]
All (hyper-) parameter settings for the algorithms/methods used in experiments have been reported, along with the rationale or method for determining them. [yes/partially/no/NA]
The number and range of (hyper-) parameter settings explored prior to conducting final experiments have been indicated, along with the effort spent on (hyper-) parameter optimisation. [yes/partially/no/NA]
Appropriately chosen statistical hypothesis tests are used to establish statistical significance in the presence of noise effects. [yes/partially/no/NA]
Articles using data sets:
Does this work rely on one or more data sets (possibly obtained from a benchmark generator or similar software artifact)? [no]
If yes, please complete the list below.
All newly introduced data sets are included in an online appendix or will be made publicly available upon publication of the paper. The online appendix follows best practices for long-term accessibility with a license that allows free usage for research purposes. [yes/partially/no/NA]
The newly introduced data set comes with a license that allows free usage for reproducibility purposes. [yes/partially/no]
The newly introduced data set comes with a license that allows free usage for research purposes in general. [yes/partially/no]
All data sets drawn from the literature or other public sources (potentially including authors’ own previously published work) are accompanied by appropriate citations. [yes/no/NA]
All data sets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. [yes/partially/no/NA]
All new data sets and data sets that are not publicly available are described in detail, including relevant statistics, the data collection process and annotation process if relevant. [yes/partially/no/NA]
All methods used for preprocessing, augmenting, batching or splitting data sets (e.g., in the context of hold-out or cross-validation) are described in detail. [yes/partially/no/NA]